You're examining some of the tiny printing on one of the newer twenty-dollar bills. A 1.5 mm tall letter appears 3 mm tall and upright when viewed through a converging lens. If the lens is held 4 cm from the bill, what is the focal length of the lens?
Answer #1

magnification

m = h'/h = 3/1.5 = 2

v/u = 2

v = 2 u

v = 2* 4 = 8 cm

now using lens formula

1/f = 1/v - 1/u

1/f = - 1/8 + 1/4

f = 8 cm

=====

Comment before rate in case any doubt, will reply for sure.. goodluck